According to the Centers for Disease Control and Prevention (CDC), 19.9 million adults (aged 18 years and older) currently have asthma, and 6.2 million children (younger than 18 years old) have asthma. This represents 8.1% of adults and 8.4% of children. The George Washington University Hospital is conducting an asthma treatment study in children and needs to enroll 720 volunteers younger than 18 years old with asthma. If 4000 children show up, estimate the probability that the number of children with asthma is at least 720.

  1. The probability is <0.0001, therefore the probability is extremely low.
  2. The probability is <0.0001, therefore the probability is extremely high.
  3. The probability is 22.27%.
  4. The probability is 8.4%.
  5. We cannot compute the probability because we do not have enough information.
Answer #1
n= 4000 p= 0.0840
here mean of distribution=μ=np= 336
and standard deviation σ=sqrt(np(1-p))= 17.5435
probability = P(X>720) = P(Z>21.89)= 1-P(Z<21.89)= 1-1= 0.0000

option A is correct The probability is <0.0001, therefore the probability is extremely low.